# Exploring lattice thermal conductivity models via interpretable deep learning to accelerate the discovery of novel materials

### v3

@@ II.1 Establishment and evaluation of LTC models
In our case, KAN was implemented using the pykan-0.0.5 library (github.com/KindXiaoming/pykan), while MLP and XGBoost were built using the PyTorch[39] and scikit-learn[40] libraries, respectively. Initially, a trial-and-error approach was employed to determine approximate ranges for the hyperparameters that might yield good performance for KAN, MLP, and XGBoost. Subsequently, we employed the Optuna[41] library to perform automated hyperparameter optimization for these models. The task of identifying optimal hyperparameters for ML models can be conceptualized as finding the optimal solution to a multivariate optimization problem. It is crucial to acknowledge that hyperparameter optimization often converges to local optima rather than the global optimum, with achieving a global optimum remaining an inherently challenging endeavor. Within the hyperparameter space, multiple local optima may exist, and no algorithm can guarantee that the solution obtained is globally optimal[42]. Nevertheless, it is generally observed that with sufficient iterations, optimization algorithms often produce models with comparable performance across different hyperparameter configurations. This suggests that while achieving global optimality remains a complex challenge, practical implementations typically deliver satisfactory results within reasonable computational effort. The hyperparameter optimization ranges are detailed in Table 1, while the final selected values are presented in Table S1.
@@ II.2 Deepening physical interpretability through sensitivity analysis
Ultimately, we retained ${{C}{\text{v}}}$, $\alpha$, $F$, and ${{K}{\text{s}}}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of LTC with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq.13), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
@@ IV.6 Implementation of DFT & GPUMD methods
For the calculation of LTC, both the phonon Boltzmann Transport Equation (BTE) method and Molecular Dynamics (MD) simulations are used. For the stable Be2C structure, specific heat and LTC are computed using ShengBTE[96]. The force constants are generated with Hiphive[97]. For other structures, specific heat is calculated using Phonopy[98], while LTC is determined through GPUMD[99]. AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ps. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,\text{\AA}^{-1}$ has been tested and found to be optimal for this study.
@@ Acknowledgements.
We would like to acknowledge the National Key R&D Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No.12122408, No.12404062, No.12302220, No.1237041184), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO. ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No. 2042023kf0109). We sincerely thank the High-Preformance Computing Center of Wuhan University for providing computational resources for our work.
@@
where $N$ represents the set of all features, $S$ is any subset of features that does not include feature $i$, $\left|S\right|$ is the number of features in set $S$, is the contribution of feature set $S$ to the model’s prediction output, and
$v\left(S\cup\left\{i\right\}\right)$
is the contribution of the feature set
$S\cup\left\{i\right\}$
, which includes feature $i$, to the model’s prediction output. The core concept of SHAP is to generate explanations for individual predictions, which can then be aggregated to provide both local and global insights into the model’s behavior[105], it can be expressed mathematically as:
@@
where
${{\phi}_{i}}\left(\mathbf{x}\right)$
is the Shapley value of the i-th variable, ${{\phi}_{0}}$ is typically set as the expected value of the training samples, $\hat{f}\left(\mathbf{x}\right)$ represents the original prediction made by the model[108].

### v4

@@ II.1 Establishment and evaluation of LTC models
In our case, KAN was implemented using the pykan-0.0.5 library (github.com/KindXiaoming/pykan), while MLP and XGBoost were built using the PyTorch[39] and scikit-learn[40] libraries, respectively. Initially, a trial-and-error approach was employed to determine approximate ranges for the hyperparameters that might yield good performance for KAN, MLP, and XGBoost. Subsequently, we employed the Optuna[41] library to perform automated hyperparameter optimization for these models. The task of identifying optimal hyperparameters for ML models can be conceptualized as finding the optimal solution to a multivariate optimization problem. It is crucial to acknowledge that hyperparameter optimization often converges to local optima rather than the global optimum, with achieving a global optimum remaining an inherently challenging endeavor. Within the hyperparameter space, multiple local optima may exist, and no algorithm can guarantee that the solution obtained is globally optimal[42]. Nevertheless, it is generally observed that with sufficient iterations, optimization algorithms often produce models with comparable performance across different hyperparameter configurations. This suggests that while achieving global optimality remains a complex challenge, practical implementations typically deliver satisfactory results within reasonable computational effort. The hyperparameter optimization ranges are detailed in Table 1, while the final selected values are presented in z1.
@@ II.2 Deepening physical interpretability through sensitivity analysis
Ultimately, we retained $C_{v}$, $\alpha$, $F$, and ${{K}{\text{s}}}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of LTC with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq.13), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
@@ IV.6 Implementation of DFT & GPUMD methods
For the calculation of LTC, both the phonon Boltzmann Transport Equation (BTE) method and Molecular Dynamics (MD) simulations are used. For the stable Be2C structure, specific heat and LTC are computed using ShengBTE[96]. The force constants are generated with Hiphive[97]. For other structures, specific heat is calculated using Phonopy[98], while LTC is determined through GPUMD[99]. AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,\text{\AA}^{-1}$ has been tested and found to be optimal for this study.
@@ Acknowledgements.
We would like to acknowledge the National Key R&D Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No.12404062, No.12474093, No.12302220, No.12374091), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO.ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No.2042023kf0109). We sincerely thank the High-Preformance Computing Center of Wuhan University for providing computational resources for our work.
@@
where $N$ represents the set of all features, $S$ is any subset of features that does not include feature $i$, $\left|S\right|$ is the number of features in set $S$, is the contribution of feature set $S$ to the model’s prediction output, and $\left(S\cup\left\{i\right\}\right)$ is the contribution of the feature set $S\cup\left\{i\right\}$, which includes feature $i$, to the model’s prediction output. The core concept of SHAP is to generate explanations for individual predictions, which can then be aggregated to provide both local and global insights into the model’s behavior[105], it can be expressed mathematically as:
@@
where ${{\phi}_{i}}\left(\mathbf{x}\right)$ is the Shapley value of the i-th variable, ${{\phi}_{0}}$ is typically set as the expected value of the training samples, $\hat{f}\left(\mathbf{x}\right)$ represents the original prediction made by the model[108].